Find the basis (a1, a2, a3) that is dual to the following basis of R3

{u1 = (1, -1, 3), u2 = (0, 1, -1), u3 = (0, 3, -2)}



1
Expert's answer
2021-02-02T12:21:19-0500

The linear functional may be expressed in the form

"a_1(x,y,z)=m_1x+m_2y+m_3z \ , \ a_2(x,y,z)=n_1x+n_2y+n_3z" ,

"a_3(x,y,z)=p_1x+p_2y+p_3z" .

By definition of dual basis ,

"a_i(u_i)=0 \ for \ i \neq j \ but \ a_i(u_i)=1 \ for \ i=j" .

We define "a_1" by setting "a_1((u_1)=1,a_1(u_2)=0,a_1(u_3)=0" , This yields

"a_1(1,-1,3)=m_1-m_2+3m_3=1 \ , \ a_1(0,1,-1)=m_2-m_3=0" ,

"a_1(0,3,-2)=3m_2-2m_3=0"

solving the system of equations yields ,

"m_1=1,m_2=0,m_3=0"

Thus "a_1(x,y,z)=x"

Now , we find "a_2" by setting , "a_2(u_1)=0,a_2(u_2)=1, a_2(u_3)=0" . This yields

"a_2(1,-1,3)=n_1-n_2+3n_3=0 \ , \ a_2(0,1,-1)=n_2-n_3=1" ,

"a_2(0,3,-2)=3n_2-2n_3=0"

Solving the system of equation yields ,

"n_1=7,n_2=-2,n_3=-3"

Thus "a_2(x,y,z)=7x-2y-3z" .

we find "a_3" by setting ,

"a_3(u_1)=0,a_3(u_2)=0,a_3(u_3)=1" , This yields

"a_3(1,-1,3)=p_1-p_2+3p_3=0 \ , \ a_3(0,1,-1)=p_2-p_3=0" ,

"a_3(0,3,-2)=3p_2-2p_3=1" ,

Solving the system of equation yields

"p_1=-2,p_2=1,p_3=1"

Thus "a_3(x,y,z)=-2x+y+z"